1.     Suppose that under the United States-Canada Air Quality Agreement, both countries agree to a combined 30 percent SO2 abatement standard. Further assume that the following SO2 abatement cost functions have been estimated by each country:

 

                                              TACUS = 500 + 1.5(AUS)2

                                              MACUS = 3 AUS

 

                                              TACCAN = 1000 + 3(ACAN)2

                                              MACCAN = 6 ACAN

 

where AUS and ACAN represent the percent of SO2 abatement achieved by the United States and Canada, respectively, and each TAC and MAC is measured in millions of dollars.

A.     If the countries implement a uniform abatement standard, find the resulting values of TAC and MAC for each nation. Based on these values, is there an economic incentive the two nations to participate in an emissions trading program? Explain. (10 marks)

B.     

 

 

 


1
Expert's answer
2021-12-08T19:54:12-0500

If uniform abatement gets implemented then each country abates half of the total tangeted abatement standard which is;"\frac{30\%}{2}=15\%"


At such standard, the TAC and MAC of each country:

US:

"TAC_{US}=500+1.5(A)^2"

"=500+1.5(15)^2"

"=837.5"


"MAC_{US}=3(A)=3(15)"

"=45"


CANADA:

"TAC_{CAN}=1000+3(A)^2"

"=1000+3(15)^2"

"=1675"


"MAC_{CAN}=6(A)=6(15)"

"=90"

As seen above, the marginal cost of both countries are not equal at uniform abatement standard therefore implying that one country incurs a higher cost than the other in the same abatement thus inefficiency occurs and produces incentives for countries to participate into emission trading program